# Distributional Lusternik–Schnirelmann category of manifolds

### v2

@@ 4. For connected sums
Let $M$ and $N$ be closed $n$-manifolds and $M\#N$ be their connected sum. An analog of the formula $\operatorname{\mathsf{cat}}(M\#N)=\max\{\operatorname{\mathsf{cat}}(M),\operatorname{\mathsf{cat}}(N)\}$ from [DS] for the classical LS-category does not hold for $\operatorname{\mathsf{dcat}}$, at least when both $\pi_{1}(M)$ and $\pi_{1}(N)$ have torsion — see [DJ, Remark 6.5] for a non-orientable example when $M=N={\mathbb{R}}P^{2}$ and Remark 5.5 for an orientable example when $M=N={\mathbb{R}}P^{3}$. It is unclear if such a general connected sum formula holds for $\operatorname{\mathsf{dcat}}$ if $\pi_{1}(N)$ is torsion-free. In this section, we provide such a partial formula when $N$ is essential and satisfies the cap property defined in Section 3.A.
Let ${\mathcal{O}}_{M}$ and ${\mathcal{O}}_{N}$ be the orientation sheaves on the closed $n$-manifolds $M$ and $N$, respectively. Let $B$ denote a closed $n$-disc common to $M$ and $N$, whose interior is deleted from $M$ and $N$ to identify the resulting boundary spheres and form the connected sum $M\#N$. If ${\mathcal{O}}_{B}$ is the orientation sheaf on $B$, then we can identify ${\mathcal{O}}_{M}$ and ${\mathcal{O}}_{N}$ along ${\mathcal{O}}_{B}$ to get the orientation sheaf ${\mathcal{O}}^{\prime}$ on $M\cup_{B}N$. Consider the quotient map $\psi:M\cup_{B}N\to(M\cup_{B}N)/B=M\vee N$. Since $B$ is contractible, $\psi$ admits a homotopy inverse $\iota:M\vee N\to M\cup_{B}N$. Then the orientation sheaf ${\mathcal{O}}$ on $M\vee N$ is obtained by pulling back ${\mathcal{O}}^{\prime}$ along $\iota$. If $q:M\#N\to M\cup_{B}N$ denotes the inclusion, then the restriction of ${\mathcal{O}}^{\prime}$ along $q$ gives the orientation sheaf ${\mathcal{O}}_{\#}$ on $M\#N$. So, in particular, the map $\psi\circ q$ takes ${\mathcal{O}}$ to ${\mathcal{O}}_{\#}$.
4.1 Theorem.
Let $M$ be a closed $n$-manifold and $N$ be a closed essential $n$-manifold such that $\pi_{1}(N)$ is torsion-free. If $N$ satisfies the cap property, then $\operatorname{\mathsf{dcat}}(M\#N)=n$. In particular, $\operatorname{\mathsf{dcat}}(M\#N)=n=\max\{\operatorname{\mathsf{dcat}}(M),\operatorname{\mathsf{dcat}}(N)\}$.
Proof.
@@
is an isomorphism for each coefficient system ${\mathcal{A}}$. Note that $(\psi\circ q)^{*}({\mathcal{O}})={\mathcal{O}}_{\#}$. Let $\kappa_{n}\in H^{n}(B\pi_{1}(N);{\mathcal{F}})$ be the primary obstruction to a section of the fibration ${\mathcal{B}}_{n}(p):P(B\pi_{1}(N))\to B\pi_{1}(N)$, see Paragraph 1 of Section 3.A for details. Then $(g\circ\phi)^{*}(\kappa_{n})\in H^{n}(M\#N;(g\circ\phi)^{*}({\mathcal{F}}))$ is the primary obstruction to the lift of $g\circ\phi$ with respect to ${\mathcal{B}}_{n}(p)$. Since $\dim(M\#N)=n$, it is the only obstruction to such a lift of $g\circ\phi$. Clearly, we have $d^{*}(g^{*}(\kappa_{n}))=\alpha\oplus g^{*}(\kappa_{n})$ for some cohomology class $\alpha$. Since $\phi=d\circ\psi\circ q$ and $(\psi\circ q)_{*}([M\#N])=[M]\oplus[N]$, we get the following set of equalities in ${\mathcal{O}}\otimes\hskip 1.42262pt(g\circ d)^{*}({\mathcal{F}})$ coefficients:
@@
Thus, $\operatorname{\mathsf{dcat}}(M\#N)=n$. Finally, we note that $\operatorname{\mathsf{dcat}}(N)=n$ due to Proposition 3.4. Therefore, $\operatorname{\mathsf{dcat}}(M\#N)=n=\max\{\operatorname{\mathsf{dcat}}(M),\operatorname{\mathsf{dcat}}(N)\}$. ∎
It follows from Theorem 4.1 and the proof of Corollary 3.5 that if $N$ is a closed essential $n$-manifold such that $B\pi_{1}(N)$ is a closed $n$-manifold, then $\operatorname{\mathsf{dcat}}(M\#N)=n$ for each closed $n$-manifold $M$. Therefore, we can directly use our symplectically aspherical manifolds from Examples 3.103.11, and 3.12 to get more examples of manifolds whose $\operatorname{\mathsf{dcat}}$ will be maximum.
@@
Let $M$ and $N$ be closed $n$-manifolds. If $N$ is aspherical, then we have $\operatorname{\mathsf{dcat}}(M\#N)=n=\max\{\operatorname{\mathsf{dcat}}(M),\operatorname{\mathsf{dcat}}(N)\}$.
@@
Since $N$ is aspherical, we can take $g$ as the identity map in the proof of Theorem 4.1. Using Poincaré duality in $N$ in ${\mathcal{O}}_{N}\otimes{\mathcal{F}}$ coefficients, we then get $n\leq\operatorname{\mathsf{dcat}}(\phi)\leq\operatorname{\mathsf{dcat}}(M\#N)\leq n$ by proceeding in the same way as above. ∎
We note that just like Corollary 3.5, Corollary 4.3 also recovers Corollary 3.6: if $N$ is a closed aspherical $n$-manifold, then $\operatorname{\mathsf{dcat}}(N)=\operatorname{\mathsf{dcat}}(S^{n}\#N)=n$. In fact, Proposition 3.4 is obtained as a special case of Theorem 4.1.
@@
For non-aspherical spaces, their rational (or real) cup lengths were typically used to determine their $\operatorname{\mathsf{dcat}}$ values in [DJ, Section 6.1]. Given a connected sum, the cup lengths of its summands can be used to estimate its $\operatorname{\mathsf{dcat}}$. Indeed, for $\mathbb{F}\in\{{\mathbb{Q}},{\mathbb{R}}\}$, we have $\max\{c\ell_{\mathbb{F}}(M),c\ell_{\mathbb{F}}(N)\}\leq\operatorname{\mathsf{dcat}}(M\#N)\leq\operatorname{\mathsf{cat}}(M\#N)$, see [J2, Remark 8.2]. For example, $n\leq\operatorname{\mathsf{dcat}}(T^{n}\#T^{n})\leq\operatorname{\mathsf{cat}}(T^{n}\#T^{n})=n$ implies $\operatorname{\mathsf{dcat}}(T^{n}\#T^{n})=n$. But from this technique alone, we cannot prove that $\operatorname{\mathsf{dcat}}(M\#N)$ is maximum unless $c\ell_{\mathbb{F}}(M)=\operatorname{\mathsf{cat}}(M\#N)$ or $c\ell_{\mathbb{F}}(N)=\operatorname{\mathsf{cat}}(M\#N)$. So, for an aspherical $n$-manifold $N$, Corollary 4.3 gives new computations $\operatorname{\mathsf{dcat}}(N\#N)$ especially in the case when $c\ell_{\mathbb{F}}(N)<\operatorname{\mathsf{cat}}(N)$ for $\mathbb{F}\in\{{\mathbb{Q}},{\mathbb{R}}\}$. We now discuss several such examples.
@@
For odd $n\geq 3$, a flat orientable $n$-manifold $\mathcal{N}$ is called a generalized Hantzsche–Wendt $n$-manifold if its holonomy group is ${\mathbb{Z}}_{2}^{n-1}$. When $n=3$, this is the well-known Hantzsche–Wendt manifold (see, for example, [Scz]), which is the union along the boundaries of twisted orientable interval-bundles over a Klein bottle (described explicitly in Section 8.B as $\mathcal{M}_{6}$). In each odd dimension $\geq 3$, $\mathcal{N}$ is a rational homology sphere due to [Scz]. Therefore, $c\ell_{{\mathbb{Q}}}(\mathcal{N})=1<n=\operatorname{\mathsf{cat}}(\mathcal{N})$. By Corollary 4.3, $\operatorname{\mathsf{dcat}}(M^{n}\#\mathcal{N})=n$, and so, $\operatorname{\mathsf{dcat}}(\mathcal{N}\#\mathcal{N})=n$ are new computations. In fact, for any $m\geq 1$, if $T^{m}$ is the $m$-torus, then $c\ell_{{\mathbb{Q}}}(\mathcal{N}\times T^{m})<n+m$. Hence, Corollary 4.3 gives a new computation $\operatorname{\mathsf{dcat}}((\mathcal{N}\times T^{m})\#(\mathcal{N}\times T^{m}))=n+m$ in each dimension $\geq 4$.
@@ 5.A. $3$-manifolds
When $\pi_{1}(M)=0$, then $\operatorname{\mathsf{dcat}}(M)\leq\operatorname{\mathsf{cat}}(M)=1\implies\operatorname{\mathsf{dcat}}(M)=1$. When $\pi_{1}(M)$ is free and non-trivial, we use some ideas from [OR, Proposition 5.1] to get the following statement for $\operatorname{\mathsf{dcat}}$.
5.1 Proposition.
If $\pi_{1}(M)\neq 0$ is free, then $\operatorname{\mathsf{dcat}}(M)=2$.
Proof.
For some $k\geq 1$, let $M=M_{1}\#\cdots\#M_{k}$ be the unique (up to permutation) decomposition of $M$ into prime $3$-manifolds $M_{i}$, see [He, Theorems 3.15 and 3.21]. Since $\pi_{1}(M)\neq 0$ is free, $\pi_{1}(M_{i})$ is free for each $1\leq i\leq k$ and there exists $j$ such that $\pi_{1}(M_{j})\neq 0$. Due to [OR, Corollary 4.6], $M_{j}$ cannot be irreducible. Therefore, $M_{j}$ must be an $S^{2}$-bundle over $S^{1}$ by [He, Lemma 3.13]. Let us write $M=N\#M_{j}$, where $N$ is the connected sum of the other prime components (if $k=1$, we take $N=S^{3}$). If $M_{j}$ is orientable, then $M_{j}=S^{2}\times S^{1}$. Therefore, $2=c\ell_{{\mathbb{Q}}}(S^{2}\times S^{1})\leq\operatorname{\mathsf{dcat}}(N\#M_{j})$. If $M_{j}$ is non-orientable, then $S^{2}\times S^{1}$ is a double cover of $M_{j}$. Hence, $N\#N\#(S^{2}\times S^{1})$ is a cover of $N\#M_{j}$. Using the rational cup length of $S^{2}\times S^{1}$ in the same way as before and Theorem 2.7, we get
@@
In any case, we have $22\leq\operatorname{\mathsf{dcat}}(M)\leq\operatorname{\mathsf{cat}}(M)=$ 2. Thus, $\operatorname{\mathsf{dcat}}(M)=2$. ∎
We will use Corollary 4.3 in proving the following statement.
5.2 Proposition.
If $\pi_{1}(M)$ is not free but torsion-free, then $\operatorname{\mathsf{dcat}}(M)=3$.
Proof.
Let $M=M_{1}\#\cdots\#M_{k}$ be the (essentially unique) decomposition of $M$ into prime $3$-manifolds $M_{i}$[He, Theorems 3.15 and 3.21]. Since $\operatorname{\mathsf{cat}}(M)=3$, due to the connected sum formula [DS], there exists $j$ such that $\operatorname{\mathsf{cat}}(M_{j})=3$. Since the LS-category of $S^{2}$-bundles over $S^{1}$ is $2$, $M_{j}$ must be irreducible by [He, Lemma 3.13]. Since $\pi_{1}(M)$ is torsion-free, so is $\pi_{1}(M_{j})$. In particular, $\pi_{1}(M_{j})$ is infinite. Thus, $M_{j}$ is not a finite quotient of $S^{3}$. This and the irreducibility of $M_{j}$ implies that $M_{j}$ is aspherical. Hence, we have $M=N\#M_{j}$, where $N$ is the sum of the other summands and $M_{j}$ is aspherical. So, we apply Corollary 4.3 to get $\operatorname{\mathsf{dcat}}(M)=3$. ∎
@@
In the above proposition, we cannot drop the torsion-free hypothesis on $\pi_{1}(M)$: if $M={\mathbb{R}}P^{3}$, then $\pi_{1}(M)={\mathbb{Z}}_{2}$ is not free but $\operatorname{\mathsf{dcat}}(M)=1$.
@@
Let $M$ be a $3$-manifold admitting the spherical geometry and let $N$ be aspherical. Since $\pi_{1}(M)$ is finite, $M\#N\in\mathscr{D}$. However, Corollary 4.3 gives $\operatorname{\mathsf{dcat}}(M\#N)=3$. So, the converse of Proposition 5.2 is not true.
Because of ${\mathbb{R}}P^{3}$ and the above examples, we see that manifolds in class $\mathscr{D}$ cannot have a constant $\operatorname{\mathsf{dcat}}$ value. If $M\in\mathscr{D}$, then while $\operatorname{\mathsf{cat}}(M)=3$, one cannot estimate $\operatorname{\mathsf{dcat}}(M)$ in general. So, we compile our findings as follows.
$\operatorname{\mathsf{dcat}}(M)=\begin{cases}1&\textup{if }\pi_{1}(M)=0\\ 2&\textup{if }\pi_{1}(M)\neq 0\text{ is free}\\ 3&\textup{if }\pi_{1}(M)\text{ is not free but torsion-free}\\ ?&\textup{if }\pi_{1}(M)\text{ has torsion}\end{cases}$
@@
  • $M$ is the non-orientable $S^{2}$-bundle over $S^{1}$. Then $\operatorname{\mathsf{dcat}}(M)=2$ by the proof of Proposition 5.1.
@@
So, if such a manifold $M$ appears in the prime decomposition of a $3$-manifold $K$, then $\operatorname{\mathsf{dcat}}(K)\geq 2$ by the cup length technique used in the proof of Proposition 5.1.
@@
5.B. $4$-manifolds with free fundamental group

### v3

@@ 4. For connected sums
Let $M$ and $N$ be closed $n$-manifolds and $M\#N$ be their connected sum. An analog of the formula
$\operatorname{\mathsf{cat}}(M\#N)=\max\{\operatorname{\mathsf{cat}}(M),\operatorname{\mathsf{cat}}(N)\}$
from [DS] for the LS-category of connected sum does not hold for $\operatorname{\mathsf{dcat}}$, at least when both $\pi_{1}(M)$ and $\pi_{1}(N)$ have torsion — see [DJ, Remark 6.5] for a non-orientable example when $M=N={\mathbb{R}}P^{2}$ and Remark 5.3 for an orientable example when $M=N={\mathbb{R}}P^{3}$. It is unclear if such a general connected sum formula holds for $\operatorname{\mathsf{dcat}}$ if $\pi_{1}(N)$ is torsion-free. In this section, we provide such a partial formula when $N$ is essential and satisfies the cap property defined in Section 3.1.
Let ${\mathcal{O}}_{M}$ and ${\mathcal{O}}_{N}$ be the orientation sheaves on the closed $n$-manifolds $M$ and $N$, respectively. Let $B$ denote a closed $n$-disc common to $M$ and $N$, whose interior is deleted from $M$ and $N$ to identify the resulting boundary spheres and form the connected sum $M\#N$. If ${\mathcal{O}}_{B}$ is the orientation sheaf on $B$, then we can identify ${\mathcal{O}}_{M}$ and ${\mathcal{O}}_{N}$ along ${\mathcal{O}}_{B}$ to get the orientation sheaf ${\mathcal{O}}^{\prime}$ on $M\cup_{B}N$. Here, $M\cup_{B}N$ denotes the union of $M$ and $N$ along the closed $n$-disc $B$. Consider the quotient map $\psi:M\cup_{B}N\to(M\cup_{B}N)/B=M\vee N$. Since $B$ is contractible, $\psi$ admits a homotopy inverse $\iota:M\vee N\to M\cup_{B}N$. Then the orientation sheaf ${\mathcal{O}}$ on $M\vee N$ is obtained by pulling back ${\mathcal{O}}^{\prime}$ along $\iota$. If $q:M\#N\to M\cup_{B}N$ denotes the inclusion, then the restriction of ${\mathcal{O}}^{\prime}$ along $q$ gives the orientation sheaf ${\mathcal{O}}_{\#}$ on $M\#N$. So, in particular, the map $\psi\circ q$ takes ${\mathcal{O}}$ to ${\mathcal{O}}_{\#}$.
Proof of Theorem B.
@@
is an isomorphism for each coefficient system ${\mathcal{A}}$. Note that $(\psi\circ q)^{*}({\mathcal{O}})={\mathcal{O}}_{\#}$. Let $\kappa_{n}\in H^{n}(B\pi_{1}(N);{\mathcal{F}})$ be the primary obstruction to a section of the fibration ${\mathcal{B}}_{n}(p):P(B\pi_{1}(N))\to B\pi_{1}(N)$, see Paragraph 1 of Section 3.1 for details. Then
$(g\circ\phi)^{*}(\kappa_{n})\in H^{n}(M\#N;(g\circ\phi)^{*}({\mathcal{F}}))$
is the primary obstruction to the lift of $g\circ\phi$ with respect to the fibration ${\mathcal{B}}_{n}(p)$. Since $\dim(M\#N)=n$, it is the only obstruction to such a lift of $g\circ\phi$. Clearly, we have $d^{*}(g^{*}(\kappa_{n}))=\alpha\oplus g^{*}(\kappa_{n})$ for some cohomology class $\alpha$. Since $\phi=d\circ\psi\circ q$ and $(\psi\circ q)_{*}([M\#N])=[M]\oplus[N]$, we get the following set of equalities in ${\mathcal{O}}\otimes\hskip 1.42262pt(g\circ d)^{*}({\mathcal{F}})$ coefficients:
@@
Thus, $\operatorname{\mathsf{dcat}}(M\#N)=n$. ∎
It follows from Theorem B and the proof of Corollary 3.4 that if $N$ is a closed essential $n$-manifold such that $B\pi_{1}(N)$ is a closed $n$-manifold, then $\operatorname{\mathsf{dcat}}(M\#N)=n$ for each closed $n$-manifold $M$. Therefore, we can directly use our symplectically aspherical manifolds from Examples 3.93.10, and 3.11 to get more examples of manifolds whose $\operatorname{\mathsf{dcat}}$ will be maximum.
@@
Let $M$ and $N$ be closed $n$-manifolds. If $N$ is aspherical, then we have $\operatorname{\mathsf{dcat}}(M\#N)=\operatorname{\mathsf{cat}}(M\#N)=n$.
@@
Since $N$ is aspherical, we can take $g$ as the identity map in the proof of Theorem B. Using Poincaré duality in $N$ in ${\mathcal{O}}_{N}\otimes{\mathcal{F}}$ coefficients, we then get $n\leq\operatorname{\mathsf{dcat}}(\phi)\leq\operatorname{\mathsf{dcat}}(M\#N)\leq n$ by proceeding in the same way as above. ∎
We note that just like Corollary 3.4, Corollary 4.2 also recovers Corollary 3.5: if $N$ is a closed aspherical $n$-manifold, then $\operatorname{\mathsf{dcat}}(N)=\operatorname{\mathsf{dcat}}(S^{n}\#N)=n$. In fact, Theorem A is obtained as a special case of Theorem B.
4.3 Corollary.
If $M$ is a closed rationally essential $3$-manifold, then $\operatorname{\mathsf{dcat}}(M)=3$.
Proof.
If $M$ is a closed rationally essential $3$-manifold, then we can write $M=P\#N$ for a closed $3$-manifold $P$ and a closed aspherical $3$-manifold $N$, see [KN, Theorem 3]. Then, Corollary 4.2 implies $\operatorname{\mathsf{dcat}}(M)=3$. ∎
We note that Corollary 4.3 is not true for general closed essential $3$-manifolds in view of the fact that $\operatorname{\mathsf{dcat}}({\mathbb{R}}P^{3})=1$.
@@
For non-aspherical spaces, their rational (or real) cup-lengths were typically used to determine their $\operatorname{\mathsf{dcat}}$ values in [DJ, Section 6.1]. Given a connected sum, the cup-lengths of its summands can be used to estimate its $\operatorname{\mathsf{dcat}}$. Indeed, for $\mathbb{F}\in\{{\mathbb{Q}},{\mathbb{R}}\}$, we have that
$\max\{c\ell_{\mathbb{F}}(M),c\ell_{\mathbb{F}}(N)\}\leq\operatorname{\mathsf{dcat}}(M\#N)\leq\operatorname{\mathsf{cat}}(M\#N),$
see [J2, Remark 8.2]. For example, $n\leq\operatorname{\mathsf{dcat}}(T^{n}\#T^{n})\leq\operatorname{\mathsf{cat}}(T^{n}\#T^{n})=n$ implies $\operatorname{\mathsf{dcat}}(T^{n}\#T^{n})=n$. But from this technique alone, we cannot prove that $\operatorname{\mathsf{dcat}}(M\#N)$ is maximum unless $c\ell_{\mathbb{F}}(M)=\operatorname{\mathsf{cat}}(M\#N)$ or $c\ell_{\mathbb{F}}(N)=\operatorname{\mathsf{cat}}(M\#N)$. So, for an aspherical $n$-manifold $N$, Corollary 4.2 gives new computations $\operatorname{\mathsf{dcat}}(N\#N)$ in the case when $c\ell_{\mathbb{F}}(N)<\operatorname{\mathsf{cat}}(N)$ for $\mathbb{F}\in\{{\mathbb{Q}},{\mathbb{R}}\}$. We now discuss several such examples.
@@
For odd $n\geq 3$, a flat orientable $n$-manifold $\mathcal{N}$ is called a generalized Hantzsche–Wendt $n$-manifold if its holonomy group is ${\mathbb{Z}}_{2}^{n-1}$. When $n=3$, this is the well-known Hantzsche–Wendt manifold (see, for example, [Scz]), which is the union along the boundaries of twisted orientable interval-bundles over a Klein bottle (described explicitly in Section 8.2 as $\mathcal{M}_{6}$). In each odd dimension $\geq 3$, $\mathcal{N}$ is a rational homology sphere due to [Scz]. Therefore, $c\ell_{{\mathbb{Q}}}(\mathcal{N})=1<n=\operatorname{\mathsf{cat}}(\mathcal{N})$. By Corollary 4.2, $\operatorname{\mathsf{dcat}}(M^{n}\#\mathcal{N})=n$, and so, $\operatorname{\mathsf{dcat}}(\mathcal{N}\#\mathcal{N})=n$ are new computations. In fact, for any $m\geq 1$, if $T^{m}$ is the $m$-torus, then $c\ell_{{\mathbb{Q}}}(\mathcal{N}\times T^{m})<n+m$. Hence, we get
$\operatorname{\mathsf{dcat}}\left(\left(\mathcal{N}\times T^{m}\right)\#\left(\mathcal{N}\times T^{m}\right)\right)=n+m$
as a new computation in each dimension $\geq 4$ by Corollary 4.2.
@@ 5.1. $3$-manifolds
Using some ideas from [OR] and Theorem A, we can compute the distributional category of any closed $3$-manifold $M$ such that $\pi_{1}(M)$ is torsion-free.
Proof of Theorem C.
First, let $M$ be simply connected. In this case, we have that $\operatorname{\mathsf{dcat}}(M)\leq\operatorname{\mathsf{cat}}(M)=1\implies\operatorname{\mathsf{dcat}}(M)=1$.
Now, let us assume that $\pi_{1}(M)\neq 0$ is free. Let $M=M_{1}\#\cdots\#M_{k}$ be the unique (up to permutation) decomposition of $M$ into prime $3$-manifolds $M_{i}$, see [He, Theorems 3.15 and 3.21]. Since $\pi_{1}(M)\neq 0$ is free, $\pi_{1}(M_{i})$ is free for each $1\leq i\leq k$ and there exists $j$ such that $\pi_{1}(M_{j})\neq 0$. Due to [OR, Corollary 4.6], $M_{j}$ cannot be irreducible. Therefore, $M_{j}$ must be an $S^{2}$-bundle over $S^{1}$ by [He, Lemma 3.13]. Let us write $M=N\#M_{j}$, where $N$ is the connected sum of the other prime components (if $k=1$, we take $N=S^{3}$). If $M_{j}$ is orientable, then $M_{j}=S^{2}\times S^{1}$. Therefore, $2=c\ell_{{\mathbb{Q}}}(S^{2}\times S^{1})\leq\operatorname{\mathsf{dcat}}(N\#M_{j})$. If $M_{j}$ is non-orientable, then $S^{2}\times S^{1}$ is a double cover of $M_{j}$. Hence, $N\#N\#(S^{2}\times S^{1})$ is a cover of $N\#M_{j}$. Using the rational cup-length of $S^{2}\times S^{1}$ in the same way as before and Theorem 2.7, we get
@@
In any case, we have $22\leq\operatorname{\mathsf{dcat}}(M)\leq\operatorname{\mathsf{cat}}(M)=$ 2. Thus, $\operatorname{\mathsf{dcat}}(M)=2$.
Finally, let $\pi_{1}(M)$ be not free but torsion-free. Let $M=M_{1}\#\cdots\#M_{k}$ be the (essentially unique) decomposition of $M$ into prime $3$-manifolds $M_{i}$. Since $\operatorname{\mathsf{cat}}(M)=3$, due to the connected sum formula [DS], there exists $j$ such that $\operatorname{\mathsf{cat}}(M_{j})=3$. Since the LS-category of $S^{2}$-bundles over $S^{1}$ is $2$, $M_{j}$ must be irreducible by [He, Lemma 3.13]. Since $\pi_{1}(M)$ is torsion-free, so is $\pi_{1}(M_{j})$. In particular, $\pi_{1}(M_{j})$ is infinite. Thus, $M_{j}$ is not a finite quotient of $S^{3}$. This and the irreducibility of $M_{j}$ implies that $M_{j}$ is aspherical. Hence, we have $M=N\#M_{j}$, where $N$ is the sum of the other summands and $M_{j}$ is aspherical. So, we apply Corollary 4.2 to get $\operatorname{\mathsf{dcat}}(M)=3$. ∎
@@
In Theorem C, we cannot drop the torsion-free hypothesis on $\pi_{1}(M)$ when $\pi_{1}(M)$ is not free: $\pi_{1}({\mathbb{R}}P^{3})={\mathbb{Z}}_{2}$ is not free but $\operatorname{\mathsf{dcat}}({\mathbb{R}}P^{3})=1$.
@@
Let $M$ be a $3$-manifold admitting the spherical geometry and let $N$ be aspherical. Since $\pi_{1}(M)$ is finite, $M\#N\in\mathscr{D}$. However, Corollary 4.2 gives $\operatorname{\mathsf{dcat}}(M\#N)=3$. So, the converse of Theorem C is not true.
Because of ${\mathbb{R}}P^{3}$ and the above examples, we see that manifolds in class $\mathscr{D}$ cannot have a constant $\operatorname{\mathsf{dcat}}$ value. If $M\in\mathscr{D}$, then while $\operatorname{\mathsf{cat}}(M)=3$, one cannot estimate $\operatorname{\mathsf{dcat}}(M)$ in general. So, Theorem C is optimal in some sense.
@@
  • $M$ is the non-orientable $S^{2}$-bundle over $S^{1}$. Then $\operatorname{\mathsf{dcat}}(M)=2$ by the proof of Theorem C.
@@
So, if such a manifold $M$ appears in the prime decomposition of a $3$-manifold $K$, then $\operatorname{\mathsf{dcat}}(K)\geq 2$ by the cup-length technique used in the proof of Theorem C.
@@
5.2. $4$-manifolds with free fundamental group